OpenStudy (anonymous):

The sum of three consecutive integers is 72. Find the integers. (2 points)22, 23, 24 25, 26, 27 23, 24, 25 24, 25, 26

OpenStudy (ksaimouli):

23,24,25 just addd three

OpenStudy (ksaimouli):

u should get 72

OpenStudy (ksaimouli):

i hop eu understood

OpenStudy (anonymous):

 s = sum =72 n = number of term = 3 a= starting number d = different between 2 number = 1 all we need is 'a' 72 =(3/2) *(2a + (3 - 1)(1) 48=2a+2 46 = 2a a=23 then it is 23,24,25